# Stronger Random Baselines for In-Context Learning

### v1

@@ Appendix A Additional details of the expected maximum random baseline
-value against random baselines.
The -value for the accuracy of a classifier against the standard random baseline is:
Line 17 follows because the binomial is a discrete distribution over integers, so if we define . By similar reasoning, we get the -value with respect to the maximum random classifier:
@@
Figure 6 shows the expected maximum accuracy for various parameter settings of , the number of examples in the evaluation dataset, , the probability of guessing a correct answer on each example (one divided by the number of choices), and , the number of evaluation set reuses. Figure 7 gives -values for the expected maximum accuracy for various parameter settings. The top row () corresponds to -values for the standard random baseline.
Different datasets experience different amounts of increase in the baseline when moving from standard to maximum baselines. Figure 9 shows this range for the tasks in the BIG-bench Lite and BIG-bench Hard suites. For example, when reusing the evaluation dataset times, three BIG-bench Lite datasets have a maximum random baseline that is 10 accuracy points above their corresponding standard random baselines. But four other datasets have an increase of less than 1 accuracy point. Using the -value corresponding to a given accuracy is an unambiguous way to compare to a random baseline across datasets.
Figure 9 shows that a single performance value can have wildly different chances of being achieved by a random baseline, depending on the dataset within a benchmark suite like BIG-bench Lite or BIG-bench Hard. For example, when evaluating BIG-bench Lite tasks (with only one evaluation of the validation set), a normalized score of 10 can correspond to a -value from near-zero to 0.5. Put another way, some tasks achieve a normalized score that is better than the maximum random baseline 99% of the time when they surpass a normalized score of 1. But it takes a normalized score of 35 for other tasks.
@@ Appendix B Evaluation details
For tasks with more than 200 examples, we sample 200 as a fixed set across all parameter combinations. We choose demonstrations from the remaining unsampled examples in the task. For the seven datasets with , there are only possible prompt demonstrations in the -shot setting. In these cases we compare to the proper maximum random baseline with instead of . For the same datasets, there is a subtlety regarding dataset size. Because the demonstrations come from the pool of available examples, the 1-shot, 2-shot, and 4-shot versions of the dataset differ by a few examples. Normally this would not make much of a difference, but the maximum random baseline depends on dataset size. For ease of visualization in Figures 3, 4, 11, 12, and 13, we plot the maximum random baseline for the dataset’s full number of examples, which is slightly weaker than each setting’s exact max random baseline (because the baseline increases with smaller ). All baseline judgments in Table 1 are with respect to each setting’s exact maximum random baseline.
@@
Appendix C Held-out set evaluations
@@
Appendix D Full results

### v2

@@ Appendix A Additional details of the expected maximum random baseline
The probability of randomly guessing a correct answer does not depend on label balance.
Consider a random classifier that guesses labels uniformly at random, independently across examples. Given a labeled example , the probability that this random guesser gets the correct label is , where is the number of possible labels. It does not depend on the dataset having balanced proportions of labels:
@@
Tail probabilities against random baselines.
The tail probability for the accuracy of a classifier against the standard random baseline is:
The final line follows because the binomial is a discrete distribution over integers, so if we define . By similar reasoning, we get the tail probability with respect to the maximum random classifier:
@@
Figure 6 shows the expected maximum accuracy for various parameter settings of , the number of examples in the evaluation dataset, , the probability of guessing a correct answer on each example (one divided by the number of choices), and , the number of random classifiers. Figure 7 gives tail probabilities for the expected maximum accuracy for various parameter settings. The top row () corresponds to tail probabilities for the standard random baseline.
Runtime.
A naive implementation of the maximum random baseline depends mainly on , the number of examples in the dataset. The most expensive settings that we consider still run in under 1 second for 10,000 examples.
@@ Appendix B Evaluation details
For tasks with more than 200 examples, we sample 200 examples as a fixed set across all parameter combinations. We choose demonstrations from the remaining unsampled examples in the task. For the seven datasets with , there are only possible prompt demonstrations in the -shot setting. In these cases we compare to the proper maximum random baseline with instead of . For the same datasets, there is a subtlety regarding dataset size. Because the demonstrations come from the pool of available examples, the 1-shot, 2-shot, and 4-shot versions of the dataset differ by a few examples. Normally this would not make much of a difference, but the maximum random baseline depends on dataset size. For ease of visualization in Figures 3, 4, 9, 10, and 11, we plot the maximum random baseline for the dataset’s full number of examples, which is slightly weaker than each setting’s exact max random baseline (because the baseline increases with smaller ). All baseline judgments in Table 1 are with respect to each setting’s exact maximum random baseline.
@@
Appendix C Results for non-quantized models
The main paper evaluates the performance of quantized models. Table 2 gives full equivalent results for non-quantized models. In this setting, 15.8% of results that exceed the standard baseline do not exceed the stronger random baseline.
| Base model | Instruction-tuned | |||||||
| 1-shot | 2-shot | 4-shot | 1-shot | 2-shot | 4-shot | |||
| BIG-bench Lite dataset | L O F | L O F | L O F | A O F | A O F | A O F | Total | |
| novel_concepts | 32 | 0 | ||||||
| known_unknowns | 46 | 4 | ||||||
| code_line_description | 60 | 0 | ||||||
| emoji_movie | 100 | 7 | ||||||
| conceptual_combinations | 103 | 0 | ||||||
| strange_stories | 174 | 1 | ||||||
| hindu_knowledge | 175 | 0 | ||||||
| bbq_lite_json | 200 | 2 | ||||||
| formal_fallacies_syllogisms_negation | 200 | 11 | ||||||
| language_identification | 200 | 7 | ||||||
| logical_deduction | 200 | 0 | ||||||
| play_dialog_same_or_different | 200 | 1 | ||||||
| strategyqa | 200 | 0 | ||||||
[…]
| vitaminc_fact_verification | 200 | 0 | ||||||
| winowhy | 200 | 0 | ||||||
| Baseline disagreements per model | 1 4 5 | 1 2 1 | 1 3 1 | 1 2 4 | 2 1 5 | 3 0 3 | ||
| Total baseline disagreements | 10 | 4 | 5 | 7 | 8 | 6 | 40 | |
| Total percentage flipped | 21% | 10% | 14% | 15% | 17% | 15% | 16% | |
Appendix D Held-out set evaluations
@@
Appendix E Full results